# Contextualized Policy Recovery: Modeling and Interpreting Medical Decisions with Adaptive Imitation Learning

### v3

@@ Abstract
Interpretable policy learning seeks to estimate intelligible decision policies from observed actions; however, existing models fall short by forcing a tradeoff between accuracy and interpretability. This tradeoff limits data-driven interpretations of human decision-making process. e.g. to audit medical decisions for biases and suboptimal practices, we require models of decision processes which provide concise descriptions of complex behaviors. Fundamentally, existing approaches are burdened by this tradeoff because they represent the underlying decision process as a universal policy, when in fact human decisions are dynamic and can change drastically with contextual information. Thus, we propose Contextualized Policy Recovery (CPR), which re-frames the problem of modeling complex decision processes as a multi-task learning problem in which complex decision policies are comprised of context-specific policies. CPR models each context-specific policy as a linear observation-to-action mapping, and generates new decision models on-demand as contexts are updated with new observations. CPR is compatible with fully offline and partially observable decision environments, and can be tailored to incorporate any recurrent black-box model or interpretable decision model. We assess CPR through studies on simulated and real data, achieving state-of-the-art performance on the canonical tasks of predicting antibiotic prescription in intensive care units ($+22\%$ AUROC vs. previous SOTA) and predicting MRI prescription for Alzheimer’s patients ($+7.7\%$ AUROC vs. previous SOTA). With this improvement in predictive performance, CPR closes the accuracy gap between interpretable and black-box methods for policy learning, allowing high-resolution exploration and analysis of context-specific decision models.
@@ 1 Introduction
Figure 1: CPR uses patient-specific treatment contexts to estimate the agent’s decision model at each timestep. Decision models are context-specific, and each action probability is an interpretable linear combination of observed features. In this way, CPR achieves exact model-based interpretability without sacrificing representational capacity.
  • We simulate a heterogeneous Markov decision process and who that CPR empiriclaly converges to the true decision model parameters, improving representation, performance, and interpretability over black-box-only models.
2 Related work
@@
Table 1: Comparison of different policy learning algorithms. All listed model classes are designed to provide interpretability beyond what a black-box recurrent model would provide: CPR provides contextualized parametric policies, POETREE (Pace et al., 2022) provides a global tree-based policy, and INTERPOLE (Hüyük et al., 2022) provides interpretations of belief states in $h$ as a summary of patient history.
@@ 3.1 Preliminaries: Contextualized Modeling
as initially described by Al-Shedivat et al. (2020). This formulation was generalized by Lengerich et al. , allowing us to model $\mathbb{P}(\theta|c)$ by any black-box model while keeping $\mathbb{P}(\mathbf{Y}|x,\theta)$ in a simple model class parametrized by $\theta$.
@@ 3.2 Contextualized Policy Recovery
To recover a policy that is both tractable and interpretable, CPR makes a practical and clinically-minded assumption that a physician places the highest importance on the most current patient information when deciding an action. While treatment history is important, this history is primarily useful for placing the patient’s current disease presentation within a context of disease progression and past treatment attempts. To represent this information hierarchy, CPR leverages contextual and historical features to generate context-specific policy models.
$P(a_{t}|x_{1},a_{1},...,x_{t-1},a_{t-1},x_{t}):=f_{\theta_{t}}(a_{t}|x_{t})$
$\theta_{t}:=g(x_{1},a_{1},...,x_{t-1},a_{t-1})$
In our experiments, $g$ is parametrized by either a vanilla RNN or LSTM (Hochreiter & Schmidhuber, 1997), $f$ is a logistic function, actions $\mathcal{A}:=\{0,1\}$ are binary, and $\ell$ is binary cross-entropy loss. Finally, CPR applies a lasso regularizer to $\theta$ to learn robust policy parameters.
4 Experiments
4.1 Medical Datasets
@@
Table 2: Action-matching performance of imitation learning algorithms. Bolded values denote the best performance of interpretable models. Open source task-agnostic models are reported as mean $\pm$ standard error for 10 bootstrap runs. INTERPOLE and POETREE, which are task-specific or closed-source, are based on prior reports: $\dagger$ reported by Hüyük et al. (2022), $\ddagger$ reported by Pace et al. (2022). NR: No values reported. $\square$: interpretable method. $\blacksquare$: black-box method.
We apply CPR to two medical datasets for canonical imitation learning tasks, ADNI and MIMIC-III. These datasets are a prime example of partially-observable decision environments in which we are forced to learn from demonstrated behavior, and where learned policies have the potential to improve clinical operations. Empirical results show that CPR significantly outperforms other interpretable baseline models, and even performing on-par with fully black-box models for both EHR datasets. Low brier scores indicate that CPR is well calibrated while achieving SOTA AUROC and AUPRC.
@@ 4.1.1 MIMIC Antibiotics
We look at 4195 patients in the intensive care unit over up to 6 timesteps extracted from the Medical Information Mart for Intensive Care III (Johnson et al., 2016) dataset and predict antibiotic prescription based on 7 observations - temperature, hematocrit, potassium, white blood cell count (WBC), blood pressure, heart rate, and creatinine. We removed the hemoglobin feature used in previous work by Pace et al. (2022) since it is highly correlated (>0.95) with the hematocrit feature.
@@
Previous work by Pace et al. (2022) and Bica et al. (2021) described a patient’s temperature and white blood cell count (WBC) as the main drivers of antibiotic treatment decisions since these are known medical criteria to counter infections (Masterton et al., 2008). Our results paint a more nuanced picture, in which heterogeneous coefficients reflect heterogeneous priorities when designing treatment plans. CPR identifies that the influence of temperature changes based on prior antibiotic prescription. For patients who have already been prescribed antibiotics, an infection has already been detected and the doctor’s decision model shifts towards mitigating the risk of possible side effects of the antibiotics treatment rather than strictly considering the benefits of the new treatment. This shift in priorities is supported by the change in the creatinine coefficient (Figure 2c). High serum creatinine can be an indicator of impaired kidney function (Gounden et al., 2023), a possible adverse effect of antibiotics (Khalili et al., 2013); as such, a high creatinine level decreases the probability of continuing antibiotic treatment. Finally, we see that increased potassium is associated with the decision to begin antibiotics (Figure 2c). Electrolyte balance (in which potassium plays a vital role) is an important factor in infectious diseases, and reduced sodium (a 1+ ion competitor with potassium) in particular is known as a marker of viral and bacterial infections (Królicka et al., 2020).
@@ 4.1.2 ADNI MRI Scans
Following prior works by Hüyük et al. (2022) and Pace et al. (2022), we apply CPR to 1605 patients from the Alzheimer’s Disease Neuroimaging Initiative (ADNI). The canonical task is to predict at each visit whether a Magnetic Resonance Image (MRI) scan is ordered for cognitive disorder diagnosis (noa, 2018). Patient observations consist of the Clinical Dementia Rating (CDR-SB) on a severity scale (normal; questionable impairment; severe dementia) (O’Bryant et al., 2008) and the MRI outcome of the previous visit falling into 4 categories: No MRI scan, below average, average and above average hippocampal volume.
@@ 4.2 Simulations
CPR incorporates both a deep learning component and a statistical modeling component to introduce a novel mechanism of interpretability, the context-specific linear policy. Our approach differs substantially from prior interpretable methods by including a deep learning component. Naturally, we wonder if CPR’s explicit linear policy representation is key to its performance and interpretability, or if accurate and robust context-specific linear policies can also be recovered from black-box policy models using post-hoc interpretation methods. To test this, we simulate a heterogeneous, action-dependent Markov decision process (MDP) and evaluate CPR versus black-box baselines on their recovery of true simulation parameters: the true action probability and the true coefficients of a context-specific linear policy (Fig. 5). While CPR explicitly generates these context-specific linear coefficients, black-box models implicitly model these coefficients as feature gradients (i.e. linear coefficients in a first-order Taylor expansion). Akin to popular post-hoc interpretability methods like LIME (Ribeiro et al., 2016), we leverage the differentiability of RNNs $\Phi(x_{t},h)\rightarrow a_{t}$ to recover the implicit context-specific linear policies $\theta$.
@@
$P(a_{t}=1|x_{t},x_{t-\tau},a_{t-\tau},t)=1/(1+\exp(-\theta\cdot x_{t}+\epsilon_{a})),\quad\quad\theta=x_{t-\tau}\cdot(2a_{t-\tau}-1)+\frac{t}{T}+\epsilon_{\theta}$
@@ 5 Discussion
In this study, we propose contextualized policies as dynamic, interpretable, and personalized linear decision models, each representing a single step in a complex treatment process. By relating individual modeling tasks through patient-specific histories and contexts we avoid the pitfalls of common personalization methods that reduce statistical power (e.g. sample splitting and subpopulation grouping). As a result, CPR matches the performance of black-box models while retaining the interpretability of linear models. Post-hoc analysis of patient-specifc models generated by CPR reveal rare covariates with outsize effects on treatment decisions, as well as extremely subtle effects in the general population. While we apply CPR in offline and partially observable environments, CPR is directly portable to online policy inference with only subtle training modifications. As a caveat, CPR makes an explicit trade-off to enable context-specific interpretability that precludes interpreting direct effects of historical features on current actions, which may make it unsuitable for tasks where the effect of historical features are direct and not mediated by a context-specific model. However, for medical imitation learning, CPR’s policy representation is well-aligned with clinical practice, where patient history contextualizes current patient symptoms and treatment options. CPR is a step toward reinforcement learning agents that explain as they think, and promises a general purpose platform for supporting and improving complex human decisions across disparate domains.
Acknowledgments
We thank the Alzheimer’s Disease Neuroimaging Initiative and Medical Information Mart for Intensive Care for data access. We would like to thank the INTERACT Program for making this collaboration possible, as well as the Begabtenstiftung Informatik for providing funding.
@@ A.1.1 ADNI
    We follow the task described by Pace et al. (2022) and Hüyük et al. (2022), taking the preprocessing code from Hüyük et al. (2022). Overall, the dataset contains patients with at least two and a median of three visits. We exclude patients without CDRSB measurements in one of their visits.
@@ A.1.2 MIMIC
    We follow the task of predicting antibiotic prescription in the intensive care unit over up to 6 timesteps as described by Pace et al. (2022) adapting the preprocessing code provided by Jarrett et al. (2021). Patient trajectories with missing values in one of the observed features or non-consecutive measurements, patients above $100$ years or below $1$ year of age, and short stays below 3 days were eliminated. We end up with 4195 patients in our dataset. Input features were standardized and the hemoglobin feature used in previous work was removed since it is highly correlated with the hematocrit feature ($>0.95$). See Table 3 for the results using all 8 input features. Since measurements are taken as averages over each day in the ICU, we use the measurements of the previous day as observations $x_{t}$ to predict $a_{t}$ ensuring that we don’t predict actions from measurements taken after the patient was treated with antibiotics earlier in the day.
    MIMIC Antibiotics
    Class Algorithm AUROC AUPRC Brier $\downarrow$
    Interpretable Logistic regression $0.60$ $0.79$ $0.20$
    INTERPOLE Hüyük et al. (2022) NR NR NR
    INTERPOLE Hüyük et al. (2022) $0.65$ NR $0.21$
    POETREE Pace et al. (2022) $0.68$ NR $0.19$
    CPR-RNN (ours) $0.79$ $0.88$ $0.15$
    CPR-LSTM (ours) $0.80$ $0.88$ $0.15$
    Black-Box RNN $0.81$ $0.89$ $0.14$
    LSTM $0.82$ $0.89$ $0.14$
    Table 3: Action-matching performance of different policy learning algorithms on the MIMIC antibiotics task using all 8 input features. Listed performance of INTERPOLE and POETREE are from prior reports of state-of-the-art performance on these canonical datasets: † reported in Hüyük et al. (2022), ‡ reported in Pace et al. (2022). NR: No values reported.

### v4

@@ Abstract
Interpretable policy learning seeks to estimate intelligible decision policies from observed actions; however, existing models force a tradeoff between accuracy and interpretability, limiting data-driven interpretations of human decision-making processes. Fundamentally, existing approaches are burdened by this tradeoff because they represent the underlying decision process as a universal policy, when in fact human decisions are dynamic and can change drastically under different contexts. Thus, we develop Contextualized Policy Recovery (CPR), which re-frames the problem of modeling complex decision processes as a multi-task learning problem, where each context poses a unique task and complex decision policies can be constructed piece-wise from many simple context-specific policies. CPR models each context-specific policy as a linear map, and generates new policy models on-demand as contexts are updated with new observations. We provide two flavors of the CPR framework: one focusing on exact local interpretability, and one retaining full global interpretability. We assess CPR through studies on simulated and real data, achieving state-of-the-art performance on predicting antibiotic prescription in intensive care units ($+22\%$ AUROC vs. previous SOTA) and predicting MRI prescription for Alzheimer’s patients ($+7.7\%$ AUROC vs. previous SOTA). With this improvement, CPR closes the accuracy gap between interpretable and black-box methods, allowing high-resolution exploration and analysis of context-specific decision models.
Machine Learning, ICML
@@ 1 Introduction
Table 1: Comparison of interpretable policy learning algorithms. All listed model classes are designed to provide interpretability beyond what a black-box recurrent model would provide: CPR provides contextualized parametric policies, POETREE (Pace et al., 2022) provides a global tree-based policy, and INTERPOLE (Hüyük et al., 2022) provides interpretations of encoded belief states as a summary of patient history. While each model is considered loosely explainable, the adaptability of CPR’s context-specific linear policies enable substantially simpler functional forms for $f(\cdot)$ than other methods, requiring no post-processing to interpret.
  • We simulate a heterogeneous Markov decision process, where CPR empirically converges to the true decision model parameters, improving representation, performance, and interpretability over black-box-only models.
2 Related work
@@ 3.1 Preliminaries: Contextualized Modeling
$y\sim\mathbb{P}(\mathbf{Y}|x,\theta),\quad\theta\sim\mathbb{P}(\theta|c)$
@@
as initially described by (Al-Shedivat et al., 2020). This formulation was generalized by (Lengerich et al., ), allowing us to model $\mathbb{P}(\theta|c)$ by any black-box model while keeping $\mathbb{P}(\mathbf{Y}|x,\theta)$ in a simple model class parametrized by $\theta$.
@@ 3.2 Contextualized Policy Recovery
$\square$ CPR-RNN global (ours) $\mathbf{0.72\pm 0.02}$ $\mathbf{0.88\pm 0.02}$ $\mathbf{0.15\pm 0.01}$ $\mathbf{0.80\pm 0.01}$ $\mathbf{0.89\pm 0.01}$ $\mathbf{0.16\pm 0.00}$
$\square$ CPR-LSTM global (ours) $\mathbf{0.72\pm 0.02}$ $\mathbf{0.88\pm 0.02}$ $\mathbf{0.15\pm 0.01}$ $\mathbf{0.79\pm 0.01}$ $\mathbf{0.89\pm 0.01}$ $\mathbf{0.16\pm 0.00}$
@@
Table 2: Action-matching performance of imitation learning algorithms. Bolded values denote the best performance of interpretable models. Open source task-agnostic models are reported as mean $\pm$ standard error for 10 bootstrap runs. INTERPOLE and POETREE, which are task-specific or closed-source, are based on prior reports: $\dagger$ reported by (Hüyük et al., 2022), $\ddagger$ reported by (Pace et al., 2022). NR: No values reported. $\square$: interpretable method. $\blacksquare$: black-box method.
To recover a policy that is both tractable and interpretable, CPR makes a practical assumption that in many real world settings, recent information has the highest importance when making a decision, and historical information contextualizes the effect of new information. In the medicine, treatment history is useful to place a patient’s current disease presentation within a context of disease progression and past treatment attempts, but decisions are always based on the current presentation. To represent this information hierarchy, CPR leverages contextual and historical features to generate context-specific policy models.
$P(a_{t}|x_{1},a_{1},...,x_{t-1},a_{t-1},x_{t}):=f_{\theta_{t}}(a_{t}|x_{t})$
$\theta_{t}=g(x_{1},a_{1},...,x_{t-1},a_{t-1})$
In our experiments, $g$ is parametrized by either a vanilla RNN or LSTM (Hochreiter & Schmidhuber, 1997), $f$ is a logistic function, actions $\mathcal{A}:=\{0,1\}$ are binary, and $\ell$ is binary cross-entropy loss. CPR applies a lasso regularizer to $\theta$ to learn robust policy parameters.
3.3 Global Interpretability
CPR synergizes black-box components with interpretable models to enable highly accurate policy models that are directly interpretable within a given context. While context-specific policies are a novel, accurate, and highly adaptable form of interpretability, there are cases where global interpretability is still necessary. To understand the exact effect of every historical feature on every action, we develop a second version of CPR which we call CPR global. In CPR global, we leverage the CPR framework to make piece-wise updates to a globally interpretable policy, which can be decomposed as a linear combination of all observed features and actions. Using the logistic policy form,
$\operatorname{logodds}(a_{t}|x_{1},a_{1},...,x_{t-1},a_{t-1},x_{t}):=\langle\theta_{t},x_{t}\rangle+\mu_{t}$
$\theta_{t}=g(x_{1},a_{1},...,x_{t-1},a_{t-1})$
$\mu_{t}=\alpha\langle\beta_{t-1},[x_{t-1},a_{t-1}]\rangle+(1-\alpha)\mu_{t-1}$
$\beta_{t-1}=h(x_{1},a_{1},...,x_{t-2})$
Where the bias term $\mu_{t}$ is updated with every new observation through a separate linear combination with $\beta$. The hyperparameter $\alpha$ weights this update against previous updates to the bias term. This form telescopes into a globally interpretable linear combination that is extended with each new observation. As a result, each observation has an exact linear effect in the context-specific global policy.
$\langle\theta_{t},x_{t}\rangle+\alpha\langle\beta_{t-1},[x_{t-1},a_{t-1}]\rangle+\alpha(1-\alpha)\langle\beta_{t-2},[x_{t-2},a_{t-2}]\rangle...$
4 Experiments
4.1 Medical Datasets
We apply CPR to two medical datasets for canonical imitation learning tasks, ADNI and MIMIC-III. These datasets are a prime example of partially-observable decision environments in which we are forced to learn from demonstrated behavior, and where learned policies have the potential to improve clinical operations. CPR significantly outperforms other interpretable baseline models, and even performing on-par with fully black-box models for both EHR datasets (Table 2). Low brier scores indicate that CPR is well calibrated while achieving SOTA AUROC and AUPRC.
@@ 4.1.1 MIMIC Antibiotics
We look at 4195 patients in the intensive care unit over up to 6 timesteps extracted from the Medical Information Mart for Intensive Care III (Johnson et al., 2016) dataset and predict antibiotic prescription based on 7 observations - temperature, hematocrit, potassium, white blood cell count (WBC), blood pressure, heart rate, and creatinine. We removed the hemoglobin feature used in previous work by (Pace et al., 2022) since it is highly correlated (>0.95) with the hematocrit feature.
@@
Previous work by (Pace et al., 2022) and (Bica et al., 2021) described a patient’s temperature and white blood cell count (WBC) as the main drivers of antibiotic treatment decisions since these are known medical criteria to counter infections (Masterton et al., 2008). Our results paint a more nuanced picture, in which heterogeneous coefficients reflect heterogeneous priorities when designing treatment plans. CPR identifies that the influence of temperature changes based on prior antibiotic prescription. For patients who have already been prescribed antibiotics, an infection has already been detected and the doctor’s decision model shifts towards mitigating the risk of possible side effects of the antibiotics treatment rather than strictly considering the benefits. This shift in priorities is supported by the change in the creatinine coefficient (Figure 1c). High serum creatinine can be an indicator of impaired kidney function (Gounden et al., 2023), a possible adverse effect of antibiotics (Khalili et al., 2013); as such, a high creatinine level decreases the probability of continuing antibiotic treatment. Finally, we see that increased potassium is associated with the decision to begin antibiotics (Figure 1c). Electrolyte balance (in which potassium plays a vital role) is an important factor in infectious diseases, and reduced sodium (a 1+ ion competitor with potassium) in particular is known as a marker of viral and bacterial infections (Królicka et al., 2020).
Figure 11 confirms that even with global interpretability, the most important factors in treatment decisions are almost always the most recent observations, confirming the sufficiency and necessity of localized context-specific policies for representing real world behaviors.
@@ 4.1.2 ADNI MRI Scans
Following prior works by (Hüyük et al., 2022) and (Pace et al., 2022), we apply CPR to 1605 patients from the Alzheimer’s Disease Neuroimaging Initiative (ADNI). The canonical task is to predict at each visit whether a Magnetic Resonance Image (MRI) scan is ordered for cognitive disorder diagnosis (noa, 2018). Patient observations consist of the Clinical Dementia Rating (CDR-SB) on a severity scale (normal; questionable impairment; severe dementia) (O’Bryant et al., 2008) and the MRI outcome of the previous visit falling into 4 categories: No MRI scan, below average, average and above average hippocampal volume.
@@ 4.2 Simulations
CPR incorporates both a deep learning component and a statistical modeling component to introduce a novel mechanism of interpretability, the context-specific linear policy. Our approach differs substantially from prior interpretable methods by including a deep learning component. Naturally, we wonder if CPR’s explicit linear policy representation is key to its performance and interpretability, or if accurate and robust context-specific linear policies can also be recovered from black-box policy models using post-hoc interpretation methods. To test this, we simulate a heterogeneous, action-dependent Markov decision process (MDP) and evaluate CPR versus black-box baselines on their recovery of simulation parameters: the true action probability and the true coefficients of a context-specific linear policy (Fig. 4). While CPR explicitly generates these context-specific linear coefficients, black-box models implicitly model these coefficients as feature gradients (i.e. linear coefficients in a first-order Taylor expansion). Akin to popular post-hoc interpretability methods like LIME (Ribeiro et al., 2016), we leverage the differentiability of RNNs $\Phi(x_{t},h)\rightarrow a_{t}$ to recover the implicit context-specific linear policies $\theta$.
@@
$P(a_{t}=1|x_{t},x_{t-\tau},a_{t-\tau},t)=1/(1+\exp(-\theta\cdot x_{t}+\epsilon_{a}))$
$\theta=x_{t-\tau}\cdot(2a_{t-\tau}-1)+\frac{t}{T}+\epsilon_{\theta}$
@@ 5 Discussion
In this study, we propose contextualized policies as dynamic, interpretable, and personalized linear decision models, each representing a single step in a complex treatment process. By relating individual modeling tasks through patient-specific histories and contexts we avoid the pitfalls of common personalization methods that reduce statistical power (e.g. sample splitting and subpopulation grouping). As a result, CPR matches the performance of black-box models while retaining the interpretability of linear models. Analysis of patient-specifc models generated by CPR reveal rare covariates with outsize effects on treatment decisions, as well as extremely subtle effects in the general population. As a framework, CPR is highly modular and extensible. It is compatible with any recurrent black-box model or interpretable decision model, and we demonstrate this by developing both locally and globally interpretable context-specific policies under the CPR framework, and applying them with common recurrent network encoders. Together, these policy models demonstrate that local context-specific policies are often sufficient to characterize and interpret real-world behaviors. While we apply CPR in offline and partially observable environments, CPR is directly portable to online policy inference with only subtle training modifications. More broadly, CPR is a step toward reinforcement learning agents that explain as they think, and promises a general purpose platform for supporting and improving complex human decisions.
Appendix A Appendix
Anonymized code is available [^1] https://anonymous.4open.science/r/contextualized_policy_recovery-A235. All model embeddings are done with UMAP (McInnes et al., ).
@@ A.1.1 ADNI
We follow the task described by (Pace et al., 2022) and (Hüyük et al., 2022), taking the preprocessing code from (Hüyük et al., 2022). Overall, the dataset contains patients with at least two and a median of three visits. We exclude patients without CDRSB measurements in one of their visits.
@@ A.1.2 MIMIC
We follow the task of predicting antibiotic prescription in the intensive care unit over up to 6 timesteps as described by (Pace et al., 2022) adapting the preprocessing code provided by (Jarrett et al., 2021). Patient trajectories with missing values in one of the observed features or non-consecutive measurements, patients above $100$ years or below $1$ year of age, and short stays below 3 days were eliminated. We end up with 4195 patients in our dataset. Input features were standardized and the hemoglobin feature used in previous work was removed since it is highly correlated with the hematocrit feature ($>0.95$). See Table 3 for the results using all 8 input features. Since measurements are taken as averages over each day in the ICU, we use the measurements of the previous day as observations $x_{t}$ to predict $a_{t}$ ensuring that we don’t predict actions from measurements taken after the patient was treated with antibiotics earlier in the day.
MIMIC Antibiotics Class Algorithm AUROC AUPRC Brier $\downarrow$ Interpretable Logistic regression $0.60$ $0.79$ $0.20$ INTERPOLE † NR NR NR INTERPOLE ‡ $0.65$ NR $0.21$ POETREE ‡ $0.68$ NR $0.19$ CPR-RNN (ours) $0.79$ $0.88$ $0.15$ CPR-LSTM (ours) $0.80$ $0.88$ $0.15$ Black-Box RNN $0.81$ $0.89$ $0.14$ LSTM $0.82$ $0.89$ $0.14$
Table 3: Action-matching performance of different policy learning algorithms on the MIMIC antibiotics task using all 8 input features. Listed performance of INTERPOLE and POETREE are from prior reports of state-of-the-art performance on these canonical datasets: † reported in (Hüyük et al., 2022), ‡ reported in (Pace et al., 2022). NR: No values reported.
ADNI MRI scans MIMIC Antibiotics Class Algorithm $\lambda$ $k$ $\lambda$ $k$ Interpretable CPR-RNN (ours) $0.0001$ $32$ $0.0001$ $32$ CPR-LSTM (ours) $0.001$ $64$ $0.0001$ $32$ Black-Box RNN - $64$ $32$ LSTM - $64$ $64$
Table 4: Hyperparameters chosen for different models.
Figure 5: CPR uses dynamic treatment contexts to generate the agent’s decision model at each timestep. Decision models are a context-specific weighted combination of observed features. With these context-specific linear decision models, CPR achieves exact model-based interpretability without sacrificing representational capacity.